OpenStudy (anonymous):

if anyone can show me how to do these it would be very helpful my mother and I are stumped =.=.........Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. and this one too Choose the equation of the line passing through the point (1, -5) and perpendicular to y = x - 10.

OpenStudy (anonymous):

For the first one, perpendicular lines have slopes that are negative inverses of each other, so you can tell from the equation that you were given the slope is -1. So a line perpendicular to that one would have a slope of 1.

OpenStudy (anonymous):

Now that you know the slope do you know what to do?

OpenStudy (anonymous):

@Jennylee.lee Hello..0!

hero (hero):

Well, if a line is perpendicular to another line, then then the product of their slopes will be -1.

hero (hero):

So for the first problem, we have y = -x + 6. The slope of this line is -1, therefore the slope, m, of the other line must be 1. We need to write the equation of the perpendicular line in the form y = mx + b. Since m = 1, we can write: y = 1x + b or simply y = x + b Next, we insert the given point (x,y) = (-4,-2) into the equation to get: -2 = -4 + b Now we isolate and solve for b: 4 - 2 = b 2 = b Therefore the equation of the perpendicular line is y = x + 2

hero (hero):

Hopefully, my steps were clear and you can try to do the other one.